# rightmargin=0in

### v2

@@ 4.2 Experimental set-up
We evaluate the following techniques for eliciting the truth or falsehood of factual statements. For the probing-based techniques, we again extract hidden states from the residual stream over the period token. For LLaMA-13B, we again use layer 13.
Logistic regression, as in Alain & Bengio (2018) but with fixed bias $b=0$.
Mass-mean probing. We use $p_{\mathrm{mm}}^{\mathrm{iid}}$ when validating on held-out IID data and $p_{\mathrm{mm}}$ otherwise.
Contrast-Consistent Search (CCS), introduced in Burns et al. (2023). CCS is an unsupervised method: given contrast pairs of statements with opposite truth values, CCS identifies a direction along which the activations of these statements are far apart. For our contrast pairs, we pair statements from cities and neg_cities, and from larger_than and smaller_than.
Logistic regression/mass-mean probing on the likely dataset. This is used to benchmark our probes against probes trained only to classify statements as being likely/unlikely text.
Calibrated few-shot prompting. Given a dataset $\mathcal{D}$, we construct a few-shot prompt which presents statements and labels from $\mathcal{D}$ to the LLM in-context; see appendix K for example prompts. We then append the remaining statements in $\mathcal{D}$ to this prompt one-at-a-time and treat the model’s predicted next token as its classification. We calibrate predictions so that half of the statements are labeled true/false; this improves performance by a few percentage points. [^7] Since performance is very sensitive to the few-shot prompt used, we perform a hyperparameter sweep for each of our LLMs, first selecting a number $n$ of shots which seems to work best, and then sampling five $n$-shot prompts and using the one which results in the highest accuracy.
Logistic regression on the validation set (oracle). This gives an upper-bound for the accuracy of a linear probe on the validation set.
@@ 4.3 Results
LLaMA-13B results are shown in figure 5; see appendix B for LLaMA-2-13B. We highight some key observations.
Figure 5: Generalization accuracy of probes trained on LLaMA-13B layer 13 residual stream activations. The $x$-axis shows the train set, and the $y$-axis shows the test set. All probes are trained on 80% of the data. When the train set and test set are the same, we evaluate on the held-out 20%. Otherwise, we evaluate on the full test set.
Generalization accuracy is high across all techniques. For instance, no matter the technique, training probes only on datasets of statements about numerical comparisons results in a probes with $92\%+$ accuracy on Spanish-English translation. The performance of the probes relative to calibrated few-shot accuracies suggest that model outputs are being influenced by features other than the truth. On average, mass-mean probes generalize slightly better than logistic regression and CCS probes.
Training on statements and their opposites improves generalization, consistent with the MCI hypothesis.
Probes trained on true/false datasets outperform probes trained on likely. While probes trained on likely are clearly better than random on some datasets, they generally perform poorly, especially for datasets where likelihood is negatively correlated (neg_cities, neg_sp_en_trans) or approximately uncorrelated (larger_than, smaller_than) with truth. This demonstrates that LLMs linearly encodes truth-relevant information beyond the plausibility of the text.
@@ 5 Causal intervention experiments
In this section we perform experiments which measure the extent to which the probing techniques of section 4 identify directions which are causally implicated in model outputs. Overall, our goal is to cause LLMs to treat false statements introduced in context as true and vice versa.
5.1 Identifying the relevant hidden states with patching
Consider the following prompt $p$:
The Spanish word ‘jirafa’ means ‘giraffe’. This statement is: TRUE […]
The Spanish word ‘aire’ means ‘silver’. This statement is: FALSE
The Spanish word ‘uno’ means ‘floor’. This statement is:
When an LLM processes this input, which hidden states contain the information which is causally relevant for the LLM’s prediction of the next token? One way to answer this question is with a patching experiment (Meng et al., 2022; Goldowsky-Dill et al., 2023). We one-at-a-time swap hidden states from the forward pass of the clean prompt $p$ into the forward pass of the “corrupted” prompt $p_{*}$, obtained from $p$ by replacing “floor” with “one” (the correct translation). For each swap, we record the probability difference $PD=P(\texttt{TRUE})-P(\texttt{FALSE})$ between output tokens. Swaps which result in a larger $PD$ indicate hidden states which are more causally implicated in the model’s decision to output TRUE or FALSE.
Figure 6 reveals three groups of causally implicated hidden states for LLaMA-13B. The final group, labeled (c), directly encodes the model’s prediction: after applying LLaMA-13B’s decoder head directly to these hidden states, the top logits belong to the tokens “true,” “True,” and “TRUE.” The first group, labeled (a), likely stores LLaMA-13B’s representation of the words “floor” or “one.” We hypothesize that in the middle region, labeled (b), the truth value of the statement is computed and stored above the token which marks the end of the clause. [^8] This motif, where “summarized” information about a clause is stored above end-of-clause signifiers, was also noted in the concurrent work of Tigges et al. (2023). See appendix B for the corresponding figure for LLaMA-2-13B.
We note that, if this hypothesis is correct, it justifies our choice of layer 13 and token position for extracting activations in the previous two sections, as this is the most downstream hidden state in group (b) and is therefore likely to carry the model’s most enriched representation of truth. In the next section, we validate this hypothesis together with the truth directions of section 4 by intervening in our LLMs’ forward pass, shifting activations in group (b) along a truth direction.
Figure 6: Difference $PD=P(\texttt{TRUE})-P(\texttt{FALSE})$ after patching residual stream activations from the corrupted prompt $p_{*}$ into the model’s forward pass when processing $p$.
5.2 Modifying hidden states by adding a truth vector
@@
$NIE=\frac{PD^{-}_{*}-PD^{-}}{PD^{+}-PD^{-}}$
where $PD^{+}$ ($PD^{-}$) denotes the average probability difference for true (false) statements with no intervention applied, and $PD^{-}_{*}$ is the average probability difference for false statements when the intervention described above is applied. If $NIE=0$ then the intervention was wholly ineffective, whereas if $NIE=1$ it indicates that the intervention induced the model to label false statements as TRUE with as much confidence as does genuine true statements.
@@
Table 2: Results of intervention experiments. The train set column indicates the datasets and probing technique (logistic regression, mass-mean probing, or CCS) which was used to identify the truth direction. Values are normalized indirect effects (NIEs) averaged over sp_en_trans.
LLaMA-13B LLaMA-2-13B
@@
cities (LR) 0.52 0.24 0.21 0.24
cities+neg_cities (LR) 0.66 0.58 0.37 0.69
cities (MM) 0.72 1.28 0.77 0.81
cities+neg_cities (MM) 0.95 1.41 0.85 0.95
cities+neg_cities (CCS) 0.70 0.96 0.49 0.84
likely (LR) 0.01 0.10 0.05 0.06
likely (MM) 0.61 0.60 0.68 0.59
[…]
Probes trained on likely have an effect, but it is much smaller than the effects from corresponding probes trained on true/false datasets. This further suggests that LLMs are not just representing the difference between probable and improbable text.
Training on statements and their negations results in directions which are more causal. This provides evidence for the MCI hypothesis of section 3.
@@ 6.1 Limitations and future work
Our work has a number of limitations. First, we focus on simple, uncontroversial statements, and therefore cannot disambiguate truth from closely related features, such as “commonly believed” or “verifiable” (Levinstein & Herrmann, 2023). Second, we only address how to identify a truth direction; we found empirically that the optimal bias for linear probes was under-determined by many of our training sets, and so we leave the problem of identifying well-generalizing biases to future work. Third, we only study two models at the same scale, both from the LLaMA family. And finally, although the evidence in sections 4 and 5 shed light on which of the hypotheses in section 3 is correct, uncertainty remains.
@@ 6.2 Conclusion
In this work we conduct a detailed investigation of the structure of LLM representations of truth. Drawing on simple visualizations, correlational evidence, and causal evidence, we find strong reason to believe that there is a “truth direction” in LLM representations. We also introduce mass-mean probing – a simple alternative to other linear probing techniques which better identifies truth directions from true/false datasets – and localize truth representations to certain hidden states.
Acknowledgments
We thank Ziming Liu and Isaac Liao for useful suggestions regarding distinguishing true text from likely text, and Wes Gurnee, Eric Michaud, and Peter Park for many helpful discussions throughout this project. We thank David Bau for useful suggestions regarding the experiments in section 5.
We also thank Oam Patel, Hadas Orgad, Sohee Yang, and Karina Nguyen for their suggestions, as well as Helena Casademunt, Max Nadeau, and Ben Edelman for giving feedback during this paper’s preparation. Plots were made with Plotly (Plotly Technologies Inc., 2015).
@@ Appendix A Scoping of truth
In this work, we consider declarative factual statements, for example “Eighty-one is larger than fifty-four” or “The city of Denver is in Vietnam.” We scope “truth” to mean the truth or falsehood of these statements; for instance the examples given have truth values of true and false, respectively. To be clear, we list here some notions of “truth” which we do not consider in this work:
@@
Moreover, the statements under consideration in this work are all simple, unambiguous, and uncontroversial. Thus, we make no attempt to disambiguate “true statements” from the following closely-related notions:
@@
  • Statements which educated people believe
On the other hand, our statements do disambiguate the notions of “true statements” and “statements which are likely to appear in training data.” For instance, given the input China is not a country in, LLaMA-13B’s top prediction for the next token is Asia, even though this completion is false. Similarly, LLaMA-13B judges the text “Eighty-one is larger than eighty-two” to be more likely than “Eighty-one is larger than sixty-four” even though the former statement is false and the latter statement is true. As shown in section 5, probes trained only on statements of likely or unlikely text fail to accurately classify true/false statements.
Appendix B Results for LLaMA-2-13B
In this section we present results for LLaMA-13B which were omitted from the main body of the text. To begin, we reproduce figure 6. The results of this experiment, shown in figure 7 governs which hidden states we train probes and perform causal interventions on.
As with LLaMA-13B, we see that causally relevant information is stored over both the final token of the statement and over the token which marks the end of the clause. Now, however, the group of hidden states which we hypothesize stores the truth value of the statement spans a slightly different range of layers: layers 8-14 instead of layers 7-13.
Thus, when extracting activations for visualizations and probing experiments, we will now extract over the final token (the end-of-clause signifier) in layer 14. The dataset visualizations and probing results for LLaMA-2-13B are shown in figures 8 and 9.
Figure 7: Results of patching experiment for LLaMA-2-13B.
Figure 8: Visualizations of LLaMA-2-13B representations of our datasets.
Figure 9: Generalization accuracy of probes trained on LLaMA-2-13B layer 14 residual stream activations.
@@ Appendix C Emergence of linear structure across layers
Figure 10: Top two principal components of representations of datasets in the LLaMA-13B residual stream at various layers.
@@
The linear structure observed in section 3 follows the following pattern: in early layers, representations are uninformative; then, in early middle layers, salient linear structure in the top few PCs rapidly emerges, with this structure emerging later for statements with a more complicated logical structure (e.g. conjunctions); finally, the linear separation becomes more salient and exits the top few PCs in later layers. See figure 10. We hypothesize that this is due to LLMs hierarchically developing understanding of their input data, before focusing on features which are most relevant to immediate next-token prediction in later layers.
Interestingly, the misalignment between cities and neg_cities and between sp_en_trans and neg_sp_en_trans also emerges over layers. This is seen in figure 11: in layer 6, representations are uninformative; then in layer 8, the NTD of cities and neg_cities appear antipodal; finally by layer 10, the NTDs have become orthogonal.
This can be interpreted in light of the MCI hypothesis. MCI would explain figure 11 as follows: in layer 8, the top PC represents a feature which is correlated with truth on cities and anti-correlated with truth on neg_cities; in layer 10, this feature remains the top PC, while a truth feature has emerged and is PC2. Since PC1 and PC2 have opposite correlations on cities and neg_cities, the datasets appear to be orthogonal.
Figure 11: Top PCs of datasets of statements and their opposites. The representations for the datasets are independently centered by subtracting off their means; without this centering there would also be a translational displacement between datasets of statements and their negations.
Appendix D Nearly all linearly-accessible truth-relevant information is in the top PCs
Figure 12: The solid lines show the validation accuracy of a linear probe trained with logistic regression on the dataset, after projecting the representations to the top $d+1$ through $d+k$ principal components. For comparison, we also show the accuracy of linear probes trained on random $k$-dimensional projections (averaged over 50 random projections).
In section 5 we saw that true and false statements linearly separate in the top PCs. We might ask how much of this separation is captured in the top PCs and how much of it remains in the remaining PCs. The answer is that nearly all of it is in the top PCs.
One way to quantify the amount of linearly accessible information in some subspace $V$ is to project our dataset $\mathcal{D}$ onto $V$ to obtain a dataset
$\mathcal{D}_{\mathrm{proj}}=\left\{(\mathrm{proj}_{V}(x),y)\right\}_{(x,y)\in\mathcal{D}}$
and record the validation accuracy of a linear probe trained with logistic regression on $\mathcal{D}$. This is shown in figure 12 for $V$ being given by the top $k+1$ through $k+d$ principal components (i.e., the top $d$ principal components, excluding the first $k$). As shown, once the top few principal components are excluded, almost no remaining linearly-accessible information remains.
Appendix E Further visualizations
Figure 13 shows PCA visualizations of all of our datsets. As shown, datasets some datasets saliently vary along features other than the true. For instance, the three clusters of statements in companies_true_false correspond to three different templates used in making the statements in that dataset. To give another example, if we were to include all comparisons between integers $x\in\{1,\dots,99\}$ in our larger_than dataset, then the top principal components would be dominated by features representing the sizes of numbers in the statements.
Figure 13: Visualizations of LLaMA-13B representations of our datasets.
In figure 14 we also visualize our datasets in the PCA bases for other datsets, expanding figure 2. We see that although our datasets do visually separate somewhat in the top PCs of the likely dataset, text liklihood does not account for all of the separation in the top PCs.
Figure 14: Visualizations of datasets in PCA bases for other datasets. All columns contain the same data and all rows are in the same basis. See figure 2.
One might ask what the top PC of the larger_than dataset is, given that it’s not truth. Figure 15 provides an interesting suggestion: it represents the absolute value of the difference between the two numbers being compared.
Figure 15: PCA visualization of larger_than. The point representing “$x$ is larger than $y$” is colored according to $x-y$.
Appendix F Mass-mean probing in terms of Mahalanobis whitening

### v3

@@ 5.2 Experimental set-up
In this section, we measure the effect that choice of training data, probing technique, and model scale has on probe accuracy.
For training data, we use one of: cities, cities + neg_cities, larger_than, larger_than + smaller_than, or likely. By comparing probes trained on cities to probes trained on cities + neg_cities, we are able to measure the effect of increasing data diversity in a particular, targeted way: namely, we mitigate the effect of linearly-represented features which have opposite-sign correlations with the truth in cities and neg_cities. As in §4, we will extract activations at the most-downstream hidden state in group (b).
Our probing techniques are logistic regression (LR), mass-mean probing (MM), and contrast-consistent search (CCS). CCS is an unsupervised method introduced in Burns et al. (2023): given contrast pairs of statements with opposite truth values, CCS identifies a direction along which the representations of these statements are far apart. For our contrast pairs, we pair statements from cities and neg_cities, and from larger_than and smaller_than.
For test sets, we use all of our (curated and uncurated) true/false datasets. Given a training set $\mathcal{D}$, we train our probe on a random 80% split of $\mathcal{D}$. Then when evaluating accuracy on a test set $\mathcal{D}^{\prime}$, we use the remaining 20% of the data if $\mathcal{D}^{\prime}=\mathcal{D}$ and the full test set otherwise. For mass-mean probing, if $\mathcal{D}=\mathcal{D}^{\prime}$, we use $p_{\mathrm{mm}}^{\mathrm{iid}}$, and we use $p_{\mathrm{mm}}$ otherwise.
Finally, we also include as baselines calibrated few-shot prompting [^7] We first sweep over a number $n$ of shots and then resample a few $n$-shot prompts to maximize performance. The word “calibrated” means we selected a threshold for $P(\texttt{TRUE})-P(\texttt{FALSE})$ such that half of the statements are labeled true; this improves performance by a few percentage points. and – as an oracle baseline – LR on the test set.
@@ 5.3 Results
Figure 5: (a) Average accuracies over all datasets aside from those used for training. (b) Accuracies of probes for varying model scales and training data, averaged over all test sets.
For each training set, probing technique, and model scale, we report the average accuracy across test sets. We expect many readers to be interested in the full results (including test set-specific accuracies), which are reported in App. D. Calibrated few-shot prompting was a surprisingly weak baseline, so we do not report it here (but see App. D).
Training on statements and their opposites improves generalization (Fig. 5(a)). When passing from cities to cities+neg_cities, this effect is largely explained by improved generalization on neg_sp_en_trans, i.e. using training data containing the word “not” improves generalization on other negated statements. On the other hand, passing from larger_than to larger_than+smaller_than also improves performance, despite both datasets being very structurally different from the rest of our datasets. As discussed in §4.1, this suggest that training on statements and their opposites mitigates the effect certain types of non-truth features have on the probe direction.
Probes generalize better for larger models (Fig. 5). While it is unsurprising that larger models are themselves better at labeling statements as true or false, it is not obvious that linear probes trained on larger models should also generalize better. Nevertheless, for LLaMA-2-13B and 70B, generalization is generally high; for example, no matter which probing technique is used, we find that probes trained on larger_than + smaller_than get $>95\%$ accuracy on sp_en_trans. This corroborates our discussion in §4.1, in which we suggested that larger models linearly represent more general concepts concepts, like truth, which capture shared aspects of diverse inputs.
Mass-mean probes generalize about as well as other probing techniques for larger models (Fig. 5(b)). While MM underperforms LR and CCS for LLaMA-2-7B, we find for larger models performance comparable to that of other probing techniques. Further, we will see in §6 that the directions identified by MM are more causally implicated in model outputs.
[…]
@@ 6 Causal intervention experiments
In §5 we measured the quality of linear probes in terms of their classification accuracy, both in- and out-of-distribution. In this section, we perform experiments which measure the extent to which these probes identify directions which are causally implicated in model outputs Finlayson et al. (2021); Geva et al. (2023); Geiger et al. (2021). To do this, we will intervene in our model’s computation by shifting the activations in group (b) (identified in §3) along the directions identified by our linear probes. Our goal is to cause LLMs to treat false statements appearing in context as true and vice versa. Crucially—and in contrast to prior work (Li et al., 2023b)—we evaluate our interventions on OOD inputs.
Table 2: NIEs for intervention experiments, averaged over statements from sp_en_trans.
LLaMA-2-13B LLaMA-2-70B
@@
cities LR .13 .19 .55 .99
MM .77 .90 .58 .89
cities+ neg_cities LR .33 .52 .61 1.00
MM .85 .97 .81 .95
CCS .31 .73 .55 .96
larger_than LR .28 .27 .61 .96
MM .71 .79 .67 1.01
larger_than+ smaller_than LR .07 .13 .54 1.02
MM .26 .53 .66 1.03
CCS .08 .17 .57 1.02
likely LR .05 .08 .18 .46
MM .70 .54 .68 .27
6.1 Experimental set-up
Let $p$ be a linear probe trained on a true/false dataset $\mathcal{D}$. Let ${\bm{\theta}}$ be the probe direction, normalized so that $p(\mu^{-}+{\bm{\theta}})=p(\mu^{+})$ where $\mu^{+}$ and $\mu^{-}$ are the mean representations of the true and false statements in $\mathcal{D}$, respectively; in other words, we normalize ${\bm{\theta}}$ so that from the perspective of the probe $p$, adding ${\bm{\theta}}$ turns the average false statement into the average true statement. If our model encodes the truth value of statements along the direction ${\bm{\theta}}$, we would expect that replacing the representation ${\bm{x}}$ of a false statement $s$ with ${\bm{x}}+{\bm{\theta}}$ would cause the model to produce outputs consistent with $s$ being a true statement.
We use inputs of the form
The Spanish word ‘fruta’ means ‘goat’. This statement is: FALSE
The Spanish word ‘carne’ means ‘meat’. This statement is: TRUE
s. This statement is:
where s varies over sp_en_trans statements. Then for each of the probes of §5 we record:
  • $PD^{+}$ and $PD^{-}$, the average probability differences $P(\texttt{TRUE})-P(\texttt{FALSE})$ for $s$ varying over true statements or false statements in sp_en_trans, respectively,
  • $PD^{+}_{*}$ and $PD^{-}_{*}$, the average probability differences where $s$ varies over true (resp. false) statements but the probe direction ${\bm{\theta}}$ is subtracted (resp. added) to each group (b) hidden state.
Finally, we report the normalized indirect effects (NIEs)
$\frac{PD_{*}^{-}-PD^{-}}{PD^{+}-PD^{-}}\quad\text{or}\quad\frac{PD_{*}^{+}-PD^{+}}{PD^{-}-PD^{+}}$
@@ 6.1 Experimental set-up
6.2 Results
Results are shown in table 2. We summarize our main takeaways.
Mass-mean probe directions are highly causal, with MM outperforming LR and CCS in 7/8 experimental conditions, often substantially. This is true despite LR, MM, and CCS probes all have very similar sp_en_trans classification accuracies.
Training on datasets and their opposites helps for cities but not for larger_than. This is surprising, considering that probes trained on larger_than + smaller_than are more accurate on sp_en_trans than probes trained on larger_than alone (see App. D), and indicates that there is more to be understood about how training on datasets and their opposites affects truth probes.
@@ 7.1 Limitations and future work
Our work has a number of limitations. First, we focus on simple, uncontroversial statements, and therefore cannot disambiguate truth from closely related features, such as “commonly believed” or “verifiable” (Levinstein & Herrmann, 2023). Second, we study only models in the LLaMA-2 family, so it is possible that some of our results do not apply for all LLMs.
This work also raises several questions which we were unable to answer here. For instance, why were interventions with mass-mean probe directions extracted from the likely dataset so effective, despite these probes not themselves being accurate at classifying true/false statements? And why did mass-mean probing with the cities + neg_cities training data perform poorly poorly for the 70B model, despite mass-mean probing with larger_than + smaller_than performing well?
@@ 7.2 Conclusion
In this work we conduct a detailed investigation of the structure of LLM representations of truth. Drawing on simple visualizations, probing experiments, and causal evidence, we find evidence that at scale, LLMs compute and linearly represent the truth of true/false statements. We also localize truth representations to certain hidden states and introduce mass-mean probing, a simple alternative to other linear probing techniques which better identifies truth directions from true/false datasets.
Acknowledgements
We thank Ziming Liu and Isaac Liao for useful suggestions regarding distinguishing true text from likely text, and Wes Gurnee, Eric Michaud, and Peter Park for many helpful discussions throughout this project. We thank David Bau for useful suggestions regarding the experiments in Sec. 6. Thanks also to Nora Belrose for discussion about the connection between difference-in-mean probing and linear erasure.
We also thank Oam Patel, Hadas Orgad, Sohee Yang, and Karina Nguyen for their suggestions, as well as Helena Casademunt, Max Nadeau, and Ben Edelman for giving feedback during this paper’s preparation. Plots were made with Plotly (Plotly Technologies Inc., 2015). Thanks to Kevin Ro Wang for catching a typo in the statement of Thm. F.1.
@@ Appendix A Scoping of truth
In this work, we consider declarative factual statements, for example “Eighty-one is larger than fifty-four” or “The city of Denver is in Vietnam.” We scope “truth” to mean factuality, i.e. the truth or falsehood of these statements; for instance the examples given have truth values of true and false, respectively. To be clear, we list here some notions of “truth” which we do not consider in this work:
@@
Moreover, the statements under consideration in this work are all simple, unambiguous, and uncontroversial. Thus, we make no attempt to disambiguate “true statements” from closely-related notions like:
@@
  • Statements which educated people believe.
On the other hand, our statements do disambiguate the notions of “true statements” and “statements which are likely to appear in training data”; See our discussion at the end of §2.
Appendix B Full patching results
Fig. 6 shows full patching results. We see that both LLaMA-2-7B and LLaMA-2-13B display the “summarization” behavior in which information relevant to the full statement is represented over the end-of-sentence punctuation token. On the other hand, LLaMA-2-70B displays this behavior in a context-dependent way – we see it for cities but not for sp_en_trans.
Figure 6: Full patching results across all three model sizes and inputs. Results are for patching false inputs (shown) to true by changing the first token shown on the left. Numbers in parentheses are the index of the token in the full (few-shot) prompt.
@@ Appendix C Emergence of linear structure across layers
Figure 7: Projections of LLaMA-2-13B representations of datasets onto their top two PCs, across various layers.
@@
The linear structure observed in §4 follows the following pattern: in early layers, representations are uninformative; then, in early middle layers, salient linear structure in the top few PCs rapidly emerges, with this structure emerging later for statements with a more complicated logical structure (e.g. conjunctions). This is shown for LLaMA-2-13B in Fig. 7. We hypothesize that this is due to LLMs hierarchically developing understanding of their input data, progressing from surface level features to more abstract concepts.
The misalignment in Fig. 3(c) also has an interesting dependence on layer. In Fig. 8 we visualize LLaMA-2-13B representations of cities and neg_cities at various layers. In early layers (left) we see antipodal alignment as in Fig. 3(b, center). As we progress through layers, we see the axes of separation rotate to lie orthogonally, until they eventually align.
One interpretation of this is that in early layers, the model computed and linearly represented some feature (like “close association”) which correlates with truth on both cities and neg_cities but with opposite signs. In later layers, the model computed and promoted to greater salience a more abstract concept which correlates with truth across both datasets.
Figure 8: PCA visualizations of LLaMA-2-13B representations of cities and neg_cities at various layers.
Appendix D Full generalization results
Here we present the full generalization results for probes trained on LLaMA-2-70B (Fig. 9), 13B (Fig. 10), and 7B (Fig. 11). The horizontal axis shows the training data for the probe and the vertical axis shows the test set.
Figure 9: Generalization results for LLaMA-2-70B.
Figure 10: Generalization results for LLaMA-2-13B.
Figure 11: Generalization results for LLaMA-2-7B.
Appendix E Mass-mean probing in terms of Mahalanobis whitening